OpenStudy (anonymous):

solve for x log x=log(x-2)+log2

OpenStudy (anonymous):

4

OpenStudy (jamesj):

Notice first that log(x-2) + log 2 = log (2(x-2)) = log(2x - 4) Hence the LHS = RHS if and only if x = 2x - 4, or 4 = x. And indeed when x = 4, log x = log 4 = 2.log 2 = log (2) + log (2) = log(x-2) + log 2.